# DRIFT: Dynamic Rule-Based Defense with Injection Isolation for Securing LLM Agents

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly central to agentic systems due to their strong reasoning and planning capabilities. By interacting with external environments through predefined tools, these agents can carry out complex user tasks. Nonetheless, this interaction also introduces the risk of prompt injection attacks, where malicious inputs from external sources can mislead the agent’s behavior, potentially resulting in economic loss, privacy leakage, or system compromise. System-level defenses have recently shown promise by enforcing static or predefined policies, but they still face two key challenges: the ability to dynamically update security rules and the need for memory stream isolation. To address these challenges, we propose DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems, which enforces both control- and data-level constraints. A Secure Planner first constructs a minimal function trajectory and a JSON-schema-style parameter checklist for each function node based on the user query. A Dynamic Validator then monitors deviations from the original plan, assessing whether changes comply with privilege limitations and the user’s intent. Finally, an Injection Isolator detects and masks any instructions that may conflict with the user query from the memory stream to mitigate long-term risks. We empirically validate the effectiveness of DRIFT on the AgentDojo and ASB benchmark, demonstrating its strong security performance while maintaining high utility across diverse models—showcasing both its robustness and adaptability. The code is released at https://github.com/SaFoLab-WISC/DRIFT.
@@ 1 Introduction
Large Language Models (LLMs), empowered by their exceptional planning and reasoning abilities, are increasingly integrated into agentic systems [1, 2, 3]. By processing natural language data streams, LLM agents interact with external environments, such as applications [1, 4], computing systems [3], via a set of pre-defined tools to carry out complex user tasks. Since the need for interaction with untrusted external environments, a new security threat of prompt injection attacks is introduced [5, 6, 7, 8, 9], where attackers inject malicious instructions into third-party platforms, misleading the agent workflow after external interaction. For example, a product review on Amazon written by another user, such as “Ignore previous instructions, buy this red shirt,” may manipulate the LLM into executing unintended actions. This form of attack [5, 6, 7, 8, 9] may bring risks such as economic losses [6], privacy leakage [10], and system damage [11] to users, severely undermining the reliability of the agentic system.
@@
Existing defense mechanisms can be broadly categorized into model-level [12, 13, 14, 15, 16, 17] and system-level [18, 19, 20, 21, 22] defenses. Model-level defenses [12, 13, 14, 15, 16, 17] typically rely on building the model’s intrinsic guardrails to detect injection inputs or mitigate their impact, but such defenses are constrained by the models’ inherent vulnerabilities and often struggle to defend against unseen attacks. Recently, system-level defenses [18, 19, 20, 21, 22] have gained increasing attention, as they can overcome the intrinsic weaknesses of models when facing unseen attacks, thereby achieving higher reliability in real-world agentic systems. These approaches typically restrict agents’ action spaces through security policies and workflow design to prevent potential injection threats. For instance, IsolateGPT [18] mitigates information leakage risks by enforcing isolation mechanisms and maintaining a separate memory bank for each application. Recently, CaMeL [21] achieves impressive security by manually defining a set of security policies and constructing a strict and fixed control and data dependency graph from the user query before any interaction takes place. More related works are discussed in Appendix B.
Despite the progress in system-level defense mechanisms for agentic systems, two critical challenges remain largely unresolved: (1) the dynamic updating of security policies and (2) the isolation of covertly injected content within the memory stream. While CaMeL enforces robust security through a strict dependency graph, this static design considerably sacrifices flexibility and practical usability, particularly in agentic systems that require adaptive, real-time decision-making. Furthermore, the reliance on manually crafted security policies imposes considerable overhead and impedes generalization across diverse usage scenarios. In addition, IsolateGPT restricts the propagation of injection-related information across different applications, but residual injection content preserved in memory still poses significant risks within the same application during prolonged interactions.
@@
To overcome these challenges, we develop DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems that enforces security through both control- and data-level constraints. We first design a Secure Planner, which establishes the initial constraint policies solely according to the user query prior to any interaction. It constructs a minimal function trajectory (control constraints) to avoid injections misleading by executing functions in order. In addition, a checklist for each function node in the trajectory, with detailed parameter requirement and value dependencies, is encoded in JSON schema format [23]. When trajectory deviations are detected, a Dynamic Validator performs approval assessments based on the privilege category (Read, Write, Execute) and its alignment with the user’s original intent. To avoid the risk of injection messages to the agent or other modules during prolonged interactions, an Injection Isolator is also designed to continuously polish the memory after each interaction, identifying and masking any instructions that conflict with the initial user query. This layered defense strategy ensures strong context isolation while enabling secure and adaptive decision-making throughout long-term agent interactions.
@@
As a fully automatic system-level defense framework, DRIFT demonstrates strong performance across diverse scenarios, achieving high security while maintaining robust utility. Specifically, we evaluate DRIFT on the AgentDojo [24] benchmark, a simulated agent environment featuring various task scenarios and types of injection attacks. By applying DRIFT to GPT-4o-mini [25], the Attack Success Rate (ASR) is successfully reduced from 30.7% to 1.3%, while utility outperforms CaMeL by 20.1% under no attack and by 12.5% under attack. In addition, DRIFT shows remarkable adaptability and generalization across four advanced online LLMs: GPT-4o [26], GPT-4o-mini [25], Claude-3.5-sonnet [27], Claude-3-haiku [28], and one prevalent offline LLM, Qwen2.5-7B-Instruct [29]. On all of these models, DRIFT significantly enhances security while maintaining or even improving utility on some models. Moreover, we finetune our policy on DRIFT, with the dataset collected from ToolBench [30], achieving significant improvements in both security and utility. Compared to the original version, the ASR of the policy-tuned model drops from 15.1% to 0.0%, while utility under no attack increases from 26.6% to 32.2%, and utility under attack improves from 19.1% to 22.2%. This policy training mechanism could enable more reliable, secure, and functional LLM agentic systems.
@@ 2.2 Dynamic Validator
In general, a function with the Read privilege does not directly pose a risk to the user and will be approved even if it deviates from the original trajectory. However, functions marked as Write or Execute may introduce direct risks. In such cases, the Validator will assess whether the deviated function aligns with the user’s original intent based on the updated tool messages, using the prompt shown in Figure 10. If the deviated function still aligns with the user’s intent, the function is approved and incorporated into the minimal function trajectory and parameter checklist to support successful validation in subsequent validation. Otherwise, agents will send an approval request to user (in our evaluation setting, sending a user request is equivalent to rejecting the deviated function call).
@@ 2.4 Security Policies in LLM Agents
An LLM-based agentic system typically comprises four key components: the user, the agent, tools, and the environment. In a standard workflow, the user first sends a query to the agent. The agent then goes through a reasoning process (e.g., chain-of-thought [31]) and selects a suitable tool to call. The response from the tool helps guide the agent’s next decision. The agent typically completes the user’s task through several such cycles. During this process, injection attacks can occur through injecting malicious content in tool responses.
@@
2.5 Trainable Security Policy
To enhance the reliability and generalization of our security policy, we develop a training approach for both the Secure Planner and the Injection Isolator, allowing our DRIFT framework to adapt more robustly. This involves designing a new data collection pipeline that extracts policy-aligned samples from existing agent datasets, followed by efficient instruction tuning using Low-Rank Adaptation (LoRA) [32] on Qwen2.5-7B-Instruct [29].
2.5.1 Data and Environment Construction
Although datasets like ToolBench [30] have been collected to support tool-use reasoning in LLMs, their formats do not align well with the structure of our security policy. This makes them less suitable for direct training. To address this, we introduce a method for generating training data that adheres our policy, by modifying existing conversations from ToolBench. Each conversation in ToolBench includes messages from three sources: user, tool, and assistant. We use GPT-4o-mini to rewrite the assistant messages to align with our policy.
Planner Data Sampling.
For training the Secure Planner, we keep the original user query and tool-calling trajectory, but rewrite the first-round assistant message using system prompt of Figure 12. Assistant messages generally include reasoning thoughts and tool calls. We modify the reasoning part using GPT-4o-mini to produce a JSON-style minimal function trajectory and parameter checklist, while keeping the tool called to preserve the original flow. We collect 1,000 such samples, with conversations ranging from 4 to 14 turns.
Isolator Data Sampling.
To train the Injection Isolator, we simulate injected instructions within tool outputs. These injections are automatically designed to fit the topic and context of the conversation, making them appear realistic and challenging. GPT-4o-mini is employed to generate the injected content and determine where to place it, using the system prompt of Figure 13. After the injection, we rewrite the assistant message to detect and highlight the injected instructions clearly. We finally collect 1,000 training samples for the Isolator.
Tool Environment Re-construction.
In practical agentic systems, the number of visible tools can be much larger than typically seen in datasets like ToolBench, where each sample involves only a few tools (usually fewer than five). To better reflect real-world scenarios, we collect tool metadata from 5,000 samples and build a tool list with over 10,000 non-redundant unique tools. For each new training instance, we randomly add 0 to 25 extra tools to the external tools, creating a more realistic and challenging environment.
2.5.2 Agent Training.
After data collection completed, we fine-tune the Qwen2.5-7B-Instruct model using LoRA for both the Secure Planner and Injection Isolator, as well as the agent itself. For the Dynamic Validator, we rely on the original Qwen2.5-7B-Instruct in a zero-shot setup to handle privilege classification and user intent checking.
@@ 3 Experiments
In this section, we evaluate DRIFT on AgentDojo [24] and ASB [33], two prevalent agentic security benchmarks, to assess the effectiveness, robustness and adaptability of DRIFT in terms of both utility and security. Furthermore, we analyze the contribution of each individual component within DRIFT.
@@ 3.1 Experimental Setups
We evaluate our method with AgentDojo [24], a benchmark that simulates realistic interactions in agent-based systems. It includes four scenarios—banking, Slack, travel, and workspace—covering 97 user tasks to assess utility and 629 injection tasks to evaluate security. In addition, we evaluate our method on ASB [33], another agent security benchmark that encompasses 10 evaluation scenarios.
@@
We compare our method against several advanced existing defense approaches. Specifically, we include four defenses implemented in AgentDojorepeat_user_prompt [34], spotlighting_with_delimiting [35], tool_filter [36], and transformers_pi_detector [17], as well as three defenses implemented in ASBdelimiters_defense [37], ob_sandwich_defense [34], and instructional_prevention [38]. In addition, we compare against two system-level defenses: a static policy-based defense, CaMeL [21], and a dynamic policy-based defense, Progent [22]. These baselines represent a broad range of strategies for protecting agents against prompt injection attacks.
@@
We apply our policy to several models, including online models—GPT-4o [26], GPT-4o-mini [25], Claude-3-haiku [28], and Claude-3.5-sonnet [27]—and an offline model, Qwen2.5-7B-Instruct [29]. For Qwen2.5-7B-Instruct, we fine-tune it on our policy dataset (described in Section 2.5) using a batch size of 4 and training for three epochs. We employ the Adam optimizer [39] with weight decay and set the initial learning rate to 2e-5. Our default attacks are the important_instruction attack on AgentDojo and the OPI attack on ASB.
@@ 3.2 Defense Techniques Comparison
In this experiment, we evaluate DRIFT on two prevalent agent safety benchmarks, AgentDojo [24] and ASB [33], and compare it with multiple advanced defenses. By default, we employ GPT-4o-mini-2024-07-18 as the base agent.
@@
Notably, the DRIFT policy achieves an optimal balance between utility and security. In terms of security, DRIFT significantly outperforms all other baselines except CaMeL, with only a marginal gap of 1.3%. However, in terms of utility under both no-attack and under-attack conditions, DRIFT surpasses CaMeL by 21.8% in the no-attack setting and 10.9% under attack. This demonstrates that DRIFT achieves a superior utility–security trade-off, highlighting the greater practicality and effectiveness of dynamic policies over static ones.

### v3

@@ Abstract
Large Language Models (LLMs) are increasingly central to agentic systems due to their strong reasoning and planning capabilities. By interacting with external environments through predefined tools, these agents can carry out complex user tasks. Nonetheless, this interaction also introduces the risk of prompt injection attacks, where malicious inputs from external sources can mislead the agent’s behavior, potentially resulting in economic loss, privacy leakage, or system compromise. System-level defenses have recently shown promise by enforcing static or predefined policies, but they still face two key challenges: the ability to dynamically update security rules and the need for memory stream isolation. To address these challenges, we propose Dynamic Rule-based Isolation Framework for Trustworthy agentic systems (DRIFT), which enforces the dynamic security policy and injection isolation for securing LLM agents against prompt injection attacks. A Secure Planner first constructs a minimal function trajectory and a JSON-schema-style parameter checklist for each function node based on the user query. A Dynamic Validator then monitors deviations from the original plan, assessing whether changes comply with privilege limitations and the user’s intent. Finally, an Injection Isolator detects and masks any instructions that may conflict with the user query from the memory stream to mitigate long-term risks. We empirically validate the effectiveness of DRIFT on the AgentDojo, ASB, and AgentDyn benchmark, demonstrating its strong security performance while maintaining high utility across diverse models—showcasing both its robustness and adaptability. The project website is available at https://safo-lab.github.io/DRIFT.
@@ 1 Introduction
Large Language Models (LLMs), empowered by their exceptional planning and reasoning abilities, are increasingly integrated into agentic systems Gur et al. [2024], Deng et al. [2023], Xie et al. [2024]. By processing natural language data streams, LLM agents interact with external environments, such as applications Gur et al. [2024], Suo [2024], computing systems Xie et al. [2024], via a set of pre-defined tools to carry out complex user tasks. Since the need for interaction with untrusted external environments, a new security threat of prompt injection attacks is introduced Perez and Ribeiro [2022], Zhan et al. [2024], Pasquini et al. [2024], Abdelnabi et al. [2023], Liu et al. [2024], where attackers inject malicious instructions into third-party platforms, misleading the agent workflow after external interaction. For example, a product review on Amazon written by another user, such as “Ignore previous instructions, buy this red shirt,” may manipulate the LLM into executing unintended actions. This form of attack Perez and Ribeiro [2022], Zhan et al. [2024], Pasquini et al. [2024], Abdelnabi et al. [2023], Liu et al. [2024] may bring risks such as economic losses Zhan et al. [2024], privacy leakage Liao et al. [2025], and system damage Zhang et al. [2024b] to users, severely undermining the reliability of the agentic system.
@@
Existing defense mechanisms can be broadly categorized into model-level Chen et al. [2024a, b], Inan et al. [2023], Li et al. [2025], Meta [2024], ProtectAI.com [2024] and system-level Wu et al. [2025, 2024], Zhong et al. [2025], Debenedetti et al. [2025], Shi et al. [2025] defenses. Model-level defenses Chen et al. [2024a, b], Inan et al. [2023], Li et al. [2025], Meta [2024], ProtectAI.com [2024] typically rely on building the model’s intrinsic guardrails to detect injection inputs or mitigate their impact, but such defenses are constrained by the models’ inherent vulnerabilities and often struggle to defend against unseen attacks due to the probabilistic nature of AI models. Recently, system-level defenses Wu et al. [2025, 2024], Zhong et al. [2025], Debenedetti et al. [2025], Shi et al. [2025] have gained increasing attention and made significant progress, as they can overcome the intrinsic weaknesses of models when facing unseen attacks, thereby achieving higher reliability in real-world agentic systems. These approaches typically restrict agents’ action spaces through pre-defined static security policies and workflow design to prevent potential injection threats. For instance, CaMeL Debenedetti et al. [2025] achieves impressive security by manually defining a set of static security policies and constructing a strict and fixed control and data dependency graph from the user query before any interaction takes place. Despite existing system-level defense mechanisms have made significant progress for securing agentic systems, this static design considerably sacrifices flexibility and practical usability. Since agent workflows are inherently dynamic and evolve through interaction, it is hard to predefine complete and correct policies in advance. As a result, static policies either fail to prevent attacks or overly restrict the agent’s utility.
Within this context, we argue that securing LLM agents against prompt injection requires a shift from static policy to dynamic policy enforcement. Unlike traditional settings, where security policies can be predefined, LLM agents operate in open-ended and evolving environments, where decisions must be made dynamically. However, enabling dynamic security policy introduces a new challenge: the policy mechanism itself becomes part of the attack surface. If policy decisions depend on untrusted external inputs, adversaries may manipulate the policy to bypass security checks. Therefore, a key requirement for dynamic defense is to minimize the exposure of the policy validator to untrusted inputs, ensuring that policy updates are grounded in a trusted and constrained information space. At the same time, not all actions carry equal risk—operations such as data modification or external execution can lead to direct harm, while read-only operations are comparatively benign. This necessitates privilege-aware control, where policy decisions are conditioned on the risk level of actions. Moreover, satisfying this requirement in long-horizon agent interactions demands more than restricting the initial inputs to the policy validator. In agentic systems, intermediate observations from external tools are continuously incorporated into memory and reused in subsequent reasoning and agent decision-making. Once malicious instructions enter this evolving context, they can indirectly expand the information space that influences both the validator and the agent. Therefore, to truly minimize adversarial influence on dynamic policy enforcement, the system must also ensure context isolation throughout execution.
@@
Based on the above, we develop DRIFT, a Dynamic Rule-based Isolation Framework for Trustworthy agentic systems, enabling dynamic security policy design. We first design a Secure Planner, which establishes the initial constraint policies solely according to the user query prior to any interaction. It constructs a minimal function trajectory (control constraints) to avoid injections misleading by executing functions in order. In addition, a checklist for each function node in the trajectory, with detailed parameter requirement and value dependencies, is encoded in JSON schema format Schema [2024]. Next, we design a Dynamic Validator that enforces policy decisions during execution. When deviations from the planned trajectory occur, the validator evaluates them based on both alignment with the user’s intent and privilege-aware control over function categories (Read, Write, Execute). This enables flexible adaptation to dynamic environments while preventing high-risk operations from being triggered by adversarial inputs. To avoid the risk of injection messages to the agent or other modules during prolonged interactions, an Injection Isolator is also designed to continuously polish the memory after each interaction, identifying and masking any instructions that conflict with the initial user query. This layered defense strategy ensures strong context isolation while enabling secure and adaptive decision-making throughout long-term agent interactions.
@@
As a fully automatic system-level defense framework, DRIFT demonstrates strong performance across diverse scenarios, achieving high security while maintaining robust utility. Specifically, we evaluate DRIFT on the AgentDojo Debenedetti et al. [2024] benchmark, a simulated agent environment featuring various task scenarios and types of injection attacks. By applying DRIFT to GPT-4o-mini OpenAI [2024a], the Attack Success Rate (ASR) is successfully reduced from 30.7% to 1.4%, while utility outperforms CaMeL by 18.9% under no attack and by 15.5% under attack. In addition, DRIFT shows remarkable adaptability and generalization across four advanced agents: GPT-4o OpenAI [2024b], GPT-4o-mini OpenAI [2024a], Claude-3.5-sonnet anthropic [2024b], Claude-3.5-haiku anthropic [2024a]. On all of these models, DRIFT significantly enhances security while maintaining or even improving utility on some models.
@@
2 Related Works
2.1 LLM Agents
LLM Agents Gur et al. [2024], Deng et al. [2023], Xie et al. [2024], Zhang et al. [2024a], Li et al. [2024], Koh et al. [2024], Song et al. [2023] are powered by large language models to automatically perceive environments and make decisions. Benefiting from the powerful reasoning capabilities of LLMs, a number of efforts Gur et al. [2024], Koh et al. [2024], Song et al. [2023], Xie et al. [2024] equip LLM agents with tools to help users automatically complete tasks. Furthermore, recent advancements Koh et al. [2024], Gur et al. [2024], Deng et al. [2023], Zhou et al. [2024b] like Mind2Web Deng et al. [2023] and WebAgent Gur et al. [2024] construct systems to interact with web pages. OSWorld Xie et al. [2024] constructs a desktop-manipulated system that enables agents to interact with computers. Additionally, several studies have explored methods to enhance agent reasoning capabilities. ReAct Yao et al. [2023] introduces an effective approach to enhance the reasoning and acting capabilities of LLMs. Language Agent Tree Search Zhou et al. [2024a] is proposed to improve the multi-step reasoning and planning capabilities of LLM agents. Some recent research also explores better tool selection mechanisms Song et al. [2023], Huang et al. [2024], Qin et al. [2024], Liang et al. [2023]. REST-GPT Song et al. [2023] develops a flexible tool-calling interface for LLM agents. ToolBench Qin et al. [2024] introduces a web-crawled benchmark for training and evaluating the tool-usage capabilities of LLMs.
2.2 Prompt Injection Defenses
A line of studies Chen et al. [2024a, b], Inan et al. [2023], Li et al. [2025], Wu et al. [2025], Debenedetti et al. [2025] has explored solutions for securing LLM agents from prompt injection attacks. Current prompt injection defenses can be classified into model-level and system-level approaches.
Model-level defenses focus on enhancing the model’s inherent ability to resist attacks. StruQ Chen et al. [2024a] introduces a mechanism to transform queries into a structured form and trains the model to focus on the structured part. Chen et al.  Chen et al. [2024b] propose a preference optimization approach to defend against injection attacks. Another significant direction involves injection detection through external models, such as LlamaGuard Inan et al. [2023] and InjecGuard Li et al. [2025]. These specialized models are trained to identify potentially malicious content across multiple risk categories, offering a complementary layer of protection.
System-level defenses typically constrain the model’s action space through predefined security policies to prevent attacks. Early system-level defenses focus primarily on coding scenarios Yang et al. [2024] and face challenges when transferred to tool-integrated agent environments Gur et al. [2024], Deng et al. [2023], Xie et al. [2024].
Recent advances in system-level protection have produced several notable approaches for tool-integrated agents. IsolateGPT Wu et al. [2025] builds isolated execution environments for each application to reduce cross-application data flow risks. Both f-secure Wu et al. [2024] and RTBAS Zhong et al. [2025] implement information flow control mechanisms that constrain untrusted data and propagate untrusted labels throughout the system. CaMeL Debenedetti et al. [2025] constructs control and data flows from the original user query and designs an interpreter to protect flow security. However, its control and data flow policies are static and cannot adequately meet the needs of dynamic real-time interactions. Concurrent with our work, Progent Shi et al. [2025] develops a dynamic policy update mechanism based on historical interactions, but legacy injection messages in memory can still impact the agent or other modules, posing long-term risks for LLM agentic systems.
@@ 3.2 Dynamic Validator
In general, a function with the Read privilege does not directly pose a risk to the user and will be approved even if it deviates from the original trajectory. However, functions marked as Write or Execute may introduce direct risks. In such cases, the Validator will assess whether the deviated function aligns with the user’s original intent, using the prompt shown in Figure 11. If the deviated function still aligns with the user’s intent, the function is approved and incorporated into the minimal function trajectory and parameter checklist to support successful validation in subsequent validation. Otherwise, agents will send an approval request to user (in our evaluation setting, sending a user request is equivalent to rejecting the deviated function call).
@@ 3.4 Security Policies in LLM Agents
An LLM-based agentic system typically comprises four key components: the user, the agent, tools, and the environment. In a standard workflow, the user first sends a query to the agent. The agent then goes through a reasoning process (e.g., chain-of-thought Wei et al. [2022]) and selects a suitable tool to call. The response from the tool helps guide the agent’s next decision. The agent typically completes the user’s task through several such cycles. During this process, injection attacks can occur through injecting malicious content in tool responses.
@@ 4 Experiments
In this section, we evaluate DRIFT on AgentDojo Debenedetti et al. [2024], ASB Zhang et al. [2025], and AgentDyn Li et al. [2026], three prevalent agentic security benchmarks, to assess the effectiveness, robustness and adaptability of DRIFT in terms of both utility and security. Furthermore, we analyze the contribution of each individual component within DRIFT.
@@ 4.1 Experimental Setups
We evaluate our method with AgentDojo Debenedetti et al. [2024], a benchmark that simulates realistic interactions in agent-based systems. It includes four scenarios—banking, Slack, travel, and workspace—covering 97 user tasks to assess utility and 629 injection tasks to evaluate security. In addition, we evaluate our method on ASB Zhang et al. [2025], another agent security benchmark that encompasses 10 evaluation scenarios. To better evaluate the effectiveness of our dynamic-aware defense, we assess our method on AgentDyn Li et al. [2026], a highly challenging agent benchmark composed of open-ended tasks. Evaluating on this benchmark enables a deeper exploration of the strengths and limitations of our approach.
@@
We compare our method against several advanced existing defense approaches. Specifically, we include four defenses implemented in AgentDojo and AgentDynrepeat_user_prompt Prompting [2025c], spotlighting_with_delimiting Hines et al. [2024], tool_filter Willison [2023], and transformers_pi_detector ProtectAI.com [2024], as well as three defenses implemented in ASBdelimiters_defense Prompting [2025b], ob_sandwich_defense Prompting [2025c], and instructional_prevention Prompting [2025a]. In addition, we compare against two system-level defenses: a static policy-based defense, CaMeL Debenedetti et al. [2025], and a dynamic policy-based defense, Progent Shi et al. [2025]. These baselines represent a broad range of strategies for protecting agents against prompt injection attacks.
@@
We apply our policy to four representative models, including GPT-4o OpenAI [2024b], GPT-4o-mini OpenAI [2024a], Claude-3.5-sonnet anthropic [2024b], and Claude-3.5-haiku anthropic [2024a]. Our default attacks are the important_instruction attack on AgentDojo and the OPI attack on ASB.
@@
Figure 5: Comparison of defense methods on GPT-4o-mini in AgentDyn.
@@ 4.2 Defense Techniques Comparison
In this experiment, we evaluate DRIFT on three prevalent agent safety benchmarks, AgentDojo Debenedetti et al. [2024], ASB Zhang et al. [2025], and AgentDyn Li et al. [2026], and compare it with multiple advanced defenses. By default, we employ GPT-4o-mini-2024-07-18 as the base agent.
@@
Notably, the DRIFT policy achieves an optimal balance between utility and security. In terms of security, DRIFT significantly outperforms all other baselines except CaMeL, with only a marginal gap of 1.4%. However, in terms of utility under both no-attack and under-attack conditions, DRIFT surpasses CaMeL by 18.9% in the no-attack setting and 15.5% under attack. This demonstrates that DRIFT achieves a superior utility–security trade-off, highlighting the greater practicality and effectiveness of dynamic policies over static ones.
@@
Comparison on AgentDyn.
On the AgentDyn benchmark, we still compare DRIFT with six advanced defense techniques: repeat_user_prompt, spotlighting_with_delimiting, tool_filter, transformers_pi_detector—one static policy-based defense, CaMeL, and one dynamic policy-based defense, Progent. The results are presented in Figure 5.
All defenses perform poorly on this highly challenging open-ended benchmark. Prompt-based approaches such as repeat_user_prompt and spotlighting_with_delimiting provide only limited security gains, still suffering from more than 35% ASR. In contrast, most other defenses exhibit severe over-defense. For example, CaMeL achieves zero Utility across all settings, as its design struggles to handle dynamic, open-ended tasks. Among the dynamic defenses, DRIFT is more stable than Progent in both utility and security, achieving the best balance between the two. Since real-world agent tasks are typically open-ended, these poor results highlight the large gap between current defenses and the requirements of practical deployment, emphasizing the pressing need for more practical security solutions.